write an equation in the slope-intercept form for a line that passes through the given pair of points. (-2,-8), (-9,-1)
Answer by checkley71(8403)

(-1+8)/(-9+2)=7/-7 OR -1 FOR THE SLOPE
NOW FOR THE Y INTERCEPT WITH A SLOPE OF -1 WE HAVE -2+2=0 FOR THE X POINT &
-8-1(2)=-10 OR (O,-10) THEN WE HAVE Y=-X-10